# Towards Principled Assessment of Tabular Data Synthesis Algorithms

### v1

@@ 5 Utility Evaluation
5.1 Existing Metrics and Limitations
@@
Machine Learning Efficacy. Machine learning efficacy has emerged as the predominant utility metric for data synthesis [32, 65, 68, 31]. Specifically, it first forms a classification (regression) problem by treating the target attribute as labels. Then, it computes the average performance of certain classifiers or regressors in a “train on synthetic data, test on real data” (TSTR) framework [30]. This metric, ideally, should reflect the synthetic data’s ability to preserve the real data distribution through test performance.
@@
Shortfalls of Machine Learning Efficacy. The computation of this metric involves the selection of machine learning models for performance evaluation. It is known that tabular data performs differently when evaluated with various machine learning models [32]. Consequently, selecting different models for comparison can lead to divergent outcomes, complicating the process of fair assessment. Furthermore, it is expected that high-quality synthetic data should closely mimic the performance behavior of real data across a spectrum of machine learning models, which is pivotal for model selection. However, machine learning efficacy, as an average-based metric, fails to capture these performance discrepancies.
5.2 Proposed Metrics: MLA and Query Error
@@
Machine Learning Affinity (MLA). The inconsistent performance of different machine learning models is caused by the distribution shift of synthetic data. It is known that machine learning models vary in their sensitivity to such shifts, leading to divergent levels of performance degradation [37]. To accurately reflect these performance fluctuations on synthetic data, we introduce an interpretable, easy-to-compute metric called machine learning affinity. It measures the relative accuracy difference between synthetic and real data under the TSTR framework. More formally we define:
Definition 3
[…]
@@
Instead of directly averaging the performance on synthetic data, we measure the relative performance discrepancy over various machine learning models. Therefore, a low affinity not only indicates the effectiveness of the synthetic data for machine learning tasks but also represents similar performance behaviors with real ones.
@@
Query Error. Range query and point query are fundamental tasks for tabular data analysis [10]. However, they are often overlooked in the evaluation of most state-of-the-art synthesizers [65, 68, 7, 60]. To accommodate both range query and point query, we follow [41] to define the query error based on the frequency of random $k$-way queries:
Definition 4
[…]
6 SynMeter
@@
This section introduces SynMeter, a modular toolkit tailored for assessing data synthesis algorithms across three dimensions: fidelity, privacy, and utility.
6.1 Tuning Objective
@@
Most synthesizers do not provide details about how to tune the hyperparameters [65, 60, 34, 41]. Instead, default settings are used for evaluations, which may not always yield optimal results, leading to subpar performance and skewed comparisons. To address this issue, we introduce the model tuning phase within our evaluation framework to facilitate the selection of hyperparameters for synthesizers. More precisely, we propose a novel and unified tuning objective, $\mathcal{L}$, to steer the tuning process:
@@
Here, $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are coefficients assigned to fidelity, machine learning affinity, and query error. These coefficients enable the selection of optimal hyperparameters by assessing the fidelity and utility of synthetic data against the validation dataset $\mathcal{D}_{\text{val}}$. In our evaluation, we set all coefficients to 1, since we find that it can significantly improve the quality of the synthetic data compared to the default configurations. For instance, this approach can boost the fidelity of TabDDPM by 13% and its utility by at least 11% (refer to Section 7.6 for detailed comparisons). Practitioners can adjust these coefficients based on the specific requirements of their applications.
6.2 Implementation
SynMeter’s design is inherently modular, offering significant flexibility to incorporate new synthesizers and datasets. As depicted in Figure 1, SynMeter comprises four modules, each module is implemented with an abstract interface for any synthesis algorithm. We envisage that SynMeter can be used for the following purposes:
  • With eight state-of-the-art synthesis algorithms implemented, SynMeter facilitates data owners to tune, train, and select different synthesizers for data publishing.
  • It serves as a benchmark for data synthesis, providing principled evaluation metrics for comparative studies.
  • The modular architecture, connected via abstract interfaces, allows for easy integration of any synthesizers and enables adaptation into other domains (e.g., genomic data synthesis), by re-implementing module-specific processing functions.
Note that while several data synthesis libraries exist [53, 49], SynMeter is the only one that provides holistic data synthesis pipelines with principled metrics.
7 Experiments
@@
  • RQ1: How do the various synthesizers perform under the principled assessment? What are the new findings?
@@
  • RQ3: How effective are our proposed metrics and evaluation framework?
Figure 1: Overview of SynMeter.
Figure 2: Average ranking comparison for six HP synthesizers. Each vertex is the average ranking of the method across 12 datasets (outer means higher rank), and each axis is the evaluation metric. “Fidelity($\mathcal{D}_{\text{train}}$)” denotes the fidelity is evaluated on training dataset, and “Fidelity($\mathcal{D}_{\text{test}}$)” means it tested on test dataset.
Figure 3: Visualization comparisons of synthetic data generated by six HP synthesizers on Bean dataset. The blue points are real data, and the orange ones are synthetic data.
7.1 Data Synthesis Models
We study a wide range of state-of-the-art data synthesizers, from statistical methods to deep generative models. We select them as they are either generally considered to perform best in practice [43, 44], widely used [65, 47], or recently emerged [32, 7, 60]. These synthesizers can be categorized into two groups: heuristic private (HP) synthesizers and DP synthesizers.
@@ 7.1 Data Synthesis Models
  • CTGAN [65] is one of the most widely used HP synthesis algorithms. It utilizes generative adversarial networks to learn the tabular data distributions. Training techniques like conditional generation and Wasserstein loss [24] are used.
  • TVAE [65] is the state-of-the-art variational autoencoder for tabular data generation, which uses mode-specific normalization to overcome the non-Gaussian and multimodal problems of continuous distribution in tabular data.
  • TabDDPM [32] is the state-of-the-art diffusion model for data synthesis. It leverages the Gaussian diffusion process and the multinomial diffusion process to model continuous and discrete distributions respectively.
  • GReaT [7] utilizes the large language model (LLM) for data synthesis. It converts records to textual representations for LLM and generates synthetic data with prompts.
@@
  • PGM [39] is the state-of-the-art DP synthesizer, which uses probabilistic graphical models [41] to learn the dependence of low-dimensional marginals for data synthesis. It won the NIST Differential Privacy Synthetic Data Challenge [43].
  • PrivSyn [70] is a non-parametric DP synthesizer, which iteratively updates the synthetic dataset to make it match the target noise marginals. This method also shows strong performance in NIST competitions [43, 44].
  • PATE-GAN [30] shares a similar architecture with CTGAN, but leverages the Private Aggregation of Teacher Ensembles (PATE) mechanism [47] to offer DP guarantees.
  • TableDiffusion [60] is a newly proposed diffusion model for tabular data synthesis, which uses Differentially Private Stochastic Gradient Descent (DP-SGD) to enforce privacy.
@@
Our goal is not to benchmark all possible synthesizers but to focus on the best-known and broad spectrum of state-of-the-art synthesizers. Hence, some synthesizers [71, 40, 34, 31] are omitted as they are considered the variants of the algorithms we are assessing. Some approaches (e.g., TabDDPM, GReaT) are nearly simultaneously introduced, limiting opportunities for an extensive comparison. Our evaluation fills this gap by providing the first comprehensive evaluation of all state-of-the-art in a standardized setting.
Figure 4: Average ranking comparison for four DP synthesizers. Each vertex is the average ranking of the method across 12 datasets (outer means higher rank), and each axis is the evaluation metric. All methods offer provable privacy guarantees (thus all rank 1st).
Figure 5: Visualization comparisons of synthetic data generated by four DP synthesizers on Bean dataset.
Table 3: List of datasets used for evaluation.
@@ 7.2 Experimental Setups
Datasets. For systematic investigation of the performance of synthesis algorithms, we consider a diverse set of 12 real-world public datasets. These datasets have various sizes, nature, attributes, and distributions. We explicitly divide datasets into training and test with a ratio of 8:2, then split 20% of the training dataset as the validation set, which is used for model tuning. All datasets are publicly available and widely used [71, 32], and the full list is detailed in Table 3.
Evaluation Details. We use proposed fidelity, privacy, and utility metrics for evaluation. We first tune the synthesizers with our proposed tuning objective using training and validation data and train them with optimal performance. Then, synthetic data are generated by generator $\mathsf{G}$ for evaluation, where we test 10 times and report the mean and standard deviation as the final score. The implementation details of each metric are as follows:
  • We use all the one-way and two-way marginals for fidelity computation and use the mean as the final score. The Wasserstein distance for univariate distributions has a closed form and can be calculated easily. For two-way marginals, we use OPT package [18] to solve the optimal problem described in Equation 1. When the cost matrix is large ($n\geq 5,000$), we randomly sample half of the elements in the probability distributions for effective calculation.
  • To estimate the membership disclosure score for privacy evaluation, we train 80 synthesis models with randomly sampled data, and use them to compute the disclosure score for all training samples.
  • We utilize eight machine learning models for utility evaluations: SVM, Logistic Regression (or Ridge Regression), Decision Tree, Random Forest, Multilayer Perceptron (MLP), XGBoost [11], CatBoost [52], and Transformers [22]. Each model is extensively tuned on real training data to ensure optimal performance (see Appendix C for details). Performance on classification and regression is evaluated by the F1 score and RMSE, respectively.
  • We randomly construct 1,000 3-way query conditions and conduct range (point) queries for both synthetic and real data to compute the query errors.
7.3 Implementation Details
@@
We use implementations from the original papers and make some small modifications to integrate them into our evaluation framework. For PGM and PrivSyn, we discretize numerical attributes into discrete ones by binning and setting the privacy budget to infinity when compared with HP synthesis algorithms. We use the same fixed random seeds for each model to improve reproducibility. To tune the hyperparameters of data synthesis algorithms and the machine learning models, we use the Optuna library [1]. The hyperparameter search space of each model is detailed in Appendix A.
7.4 Overall Evaluation (RQ1)
Overview. Figure 2 and Figure 4 report the overview ranking results for HP and DP synthesizers, respectively. We set the privacy budget $\epsilon$ to 1 for all DP synthesis algorithms. For HP synthesizers, the diffusion model (i.e., TabDDPM) exhibits superior fidelity and utility, albeit at the expense of compromising privacy. Statistical methods like PrivSyn also achieve impressive fidelity, surpassing complex deep generative models like TVAE and GReaT. Conversely, CTGAN, the most popular HP synthesizer, shows the least satisfactory results in synthetic data quality but offers better privacy protection.
In terms of DP synthesizers, statistical methods continue to perform well in terms of fidelity and utility. However, the performance of deep generative models drops significantly to ensure differential privacy. Even the strongest diffusion-based model (i.e., TableDiffusion) underperforms statistical approaches by a large margin, which starkly contrasts with its performance in the HP context, indicating a pronounced impact of privacy constraints on deep generative models.
We also visualize the synthetic and real data by using the t-SNE technique [63], as depicted in Figure 3and Figure 5. The visualizations generally reveal a considerable overlap between synthetic and real data points, demonstrating the synthesizers’ ability to generate data resembling the original. One interesting observation is that PrivSyn is the only method that does not have clear cluster patterns. This is because PrivSyn is the only non-parametric approach that directly modifies the synthetic records to match the marginal distribution, potentially resulting in less realistic records.
@@ 7.4 Overall Evaluation (RQ1)
Table 4: Fidelity evaluation of data synthesis on train data $\mathcal{D}_{\text{train}}$. The privacy budget $\epsilon$ of HP synthesizers is set as $\infty$ (the top part), and the budget for the DP synthesizers is set as $1$ (the bottom part). The best result is in bold.
@@
[…]
@@
Table 5: Privacy evaluation (membership disclosure score) of HP synthesizers ($\epsilon=\infty$). The lower score means the synthesizer can provide better empirical membership privacy protection. The best result is in bold.

### v2

@@ 5. Utility Evaluation
5.1. Existing Metrics and Limitations
@@
Machine Learning Efficacy. Machine learning efficacy has emerged as the predominant utility metric for data synthesis (Kotelnikov et al., 2023; Xu et al., 2019; Zhang et al., 2023b; Kim et al., 2022). Specifically, it first forms a classification (regression) problem by treating the target attribute as labels. Then, it computes the average performance of certain classifiers or regressors in a “train on synthetic data, test on real data” (TSTR) framework (Jordon et al., 2018). This metric, ideally, should reflect the synthetic data’s ability to preserve the real data distribution through test performance.
@@
Shortfalls of Machine Learning Efficacy. The computation of this metric involves the selection of machine learning models (we also call “evaluators”). It is known that tabular data performs differently when evaluated with various evaluators (Kotelnikov et al., 2023), and there is no dominant model that can achieve the best performance on all tabular datasets (Gorishniy et al., 2021) (shown in Section 7.5). Consequently, selecting different evaluators for comparison can lead to divergent outcomes, making fair assessment difficult. Furthermore, it is expected that high-quality synthetic data should closely mimic the performance behavior of real data across a spectrum of machine learning models, which is pivotal for tasks like model selection. However, directly using the accuracy of certain evaluators as the utility metric fails to reflect the performance behaviors of synthetic data.
5.2. Proposed Metrics: MLA and Query Error
@@
Machine Learning Affinity (MLA). The inconsistent performance of different machine learning models is caused by the distribution shift of synthetic data. It is known that machine learning models vary in their sensitivity to such shifts, leading to divergent levels of performance degradation (Lopes et al., 2021). To accurately reflect these performance fluctuations on synthetic data, we introduce an interpretable, easy-to-compute metric called machine learning affinity. It measures the relative accuracy difference between synthetic and real data under the TSTR framework. More formally we define:
Definition 0.
(Machine Learning Affinity) Let $\mathcal{E}$ be a set of candidate machine learning models (evaluators), let $e_{\mathcal{D}_{\text{train}}}$ and $e_{\mathcal{D}_{\mathcal{S}}}$ be evaluators trained on real training data $\mathcal{D}_{\text{train}}$ and synthetic data $\mathcal{S}$, $\mathcal{A}(e,\mathcal{D}_{\text{test}})$ denotes the evaluator’s accuracy (F1 score or RMSE) when performed on test dataset $\mathcal{D}_{\text{test}}$. The MLA of synthesizer $\mathsf{A}$ is:
(8) $\mathrm{MLA}(\mathsf{A})\coloneqq\underset{\begin{subarray}{c}e\in\mathcal{E}\end{subarray}}{\mathbb{E}}\big{[}\frac{\mathcal{A}(e_{\mathcal{D}_{\text{train}}},\mathcal{D}_{\text{test}})-\mathcal{A}(e_{\mathcal{S}},\mathcal{D}_{\text{test}})}{\mathcal{A}(e_{\mathcal{D}_{\text{train}}},\mathcal{D}_{\text{test}})}\big{]}.$
@@
Instead of directly averaging the test performance, MLA measures the relative performance discrepancy over various machine learning models. Therefore, the low MLA not only indicates the effectiveness of the synthetic data for machine learning tasks but also represents similar performance behaviors with real ones.
@@
Query Error. Range query and point query are fundamental tasks for tabular data analysis (Chatfied and Collins, 2013). However, they are often overlooked in the evaluation of most state-of-the-art synthesizers (Xu et al., 2019; Zhang et al., 2023b; Borisov et al., 2023; Truda, 2023). To accommodate both range query and point query, we follow (McKenna et al., 2019) to define the query error as below:
Definition 0.
(Query Error) Consider a subset of $k$ attributes $a=\{a_{1},a_{2},...,a_{k}\}$ sampled from data $\mathcal{D}$. For each attribute, if $a_{i}$ is categorical, a value $v_{i}$ is randomly chosen from its domain $\mathbb{R}(a_{i})$, which forms the basis for a point query condition; for numerical attributes, two values $s_{i}$ and $d_{i}$ from $\mathbb{R}(a_{i})$ are randomly sampled as the start and end points, to construct a range query condition. The final query $c\in\mathcal{C}$ combines $k$ sub-queries and is executed on both real and synthetic data to obtain query frequency ratios $\mu_{c}^{\mathcal{D}_{\text{test}}}$ and $\mu_{c}^{\mathcal{S}}$. The query error of synthesis algorithm $\mathsf{A}$ is defined as:
(9) $\mathrm{QueryError}(\mathsf{A})\coloneqq\underset{\begin{subarray}{c}c\in\mathcal{C}\end{subarray}}{\mathbb{E}}\big{[}||\mu_{c}^{\mathcal{D}_{\text{test}}}-\mu_{c}^{\mathcal{S}}||_{1}\big{]}.$
Implementations. We utilize eight machine learning models to compute MLA: SVM, Logistic Regression (or Ridge Regression), Decision Tree, Random Forest, Multilayer Perceptron (MLP), XGBoost (Chen and Guestrin, 2016), CatBoost (Prokhorenkova et al., 2018), and Transformers (Gorishniy et al., 2021). Each model is extensively tuned on real training data to ensure optimal performance (refer to the technical report version for details). Performance on classification and regression is evaluated by the F1 score and RMSE, respectively. For query error, we randomly construct 1,000 3-way query conditions and conduct range (point) queries for both synthetic and real data.
6. $\mathsf{SynMeter}$
@@
This section introduces $\mathsf{SynMeter}$, a modular toolkit tailored for assessing data synthesis algorithms across three dimensions: fidelity, privacy, and utility.
6.1. Tuning Objective
@@
Most synthesizers do not provide details about how to tune their hyperparameters (Xu et al., 2019; Truda, 2023; Lee et al., 2023; McKenna et al., 2019). Instead, default settings are used for evaluations, which may not always yield optimal results, leading to subpar performance and skewed comparisons. To address this issue, we introduce the model tuning phase within our evaluation framework to facilitate the selection of hyperparameters for synthesizers. More precisely, we propose a simple and unified tuning objective, $\mathcal{L}$, to steer the tuning process:
@@
Here, $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ are coefficients assigned to fidelity (i.e., Wasserstein distance), machine learning affinity, and query error. In our evaluation, we set all coefficients to 1, since we find that it can significantly improve the quality of the synthetic data compared to default configurations. We utilize Optuna (Akiba et al., 2019) for hyperparameter optimizations by minimizing $\mathcal{L}$.
Discussion. Despite the simplicity of the setup described, we observe consistent improvement when applying this tuning strategy. For instance, this approach can boost the fidelity of TabDDPM by 13% and utility by at least 11% (refer to Section 7.5 for detailed comparisons). In practice, the values of fidelity and utility metrics fall within the same scale (although they use different measurements), which alleviates the need to assign different weights for each aspect. Practitioners can adjust these coefficients based on the specific requirements of applications (e.g., higher $\alpha_{2}$ for model selection (Shoshan et al., 2023)). We leave other assignment strategies for future work.
It is worth mentioning that we do not include the privacy metric for model tuning. The reason is two-fold: (1) DP synthesizers already offer provable privacy guarantees, eliminating the need for privacy-oriented tuning. By excluding the privacy metric, we establish a more generalized tuning objective that can apply to both HP and DP synthesizers. (2) Empirically we find that incorporating MDS into the tuning objective $\mathcal{L}$ results in negligible improvements in the privacy of synthetic data (this may be because some synthesizers are inherently private, discussed in Section 7.5). Moreover, computing MDS requires training multiple models for each tuning iteration, which can significantly hinder the efficiency of the tuning phase.
6.2. Implementation
$\mathsf{SynMeter}$’s design is inherently modular, offering significant flexibility to incorporate new synthesizers and datasets. As depicted in Figure 1, $\mathsf{SynMeter}$ comprises four modules, each module is implemented with an abstract interface for any synthesis algorithm. We envisage that $\mathsf{SynMeter}$ can be used for the following purposes:
  • With eight state-of-the-art synthesis algorithms implemented, $\mathsf{SynMeter}$ facilitates data owners to tune, train, and select different synthesizers for data publishing.
  • It serves as a benchmark for data synthesis, providing systematic evaluation metrics for comparative studies.
  • The modular architecture, connected via abstract interfaces, allows for easy integration of any synthesizers and enables adaptation into other domains (e.g., genomic data synthesis (Oprisanu et al., 2022)), by re-implementing module-specific processing functions.
Figure 1. Overview of $\mathsf{SynMeter}$.
7. Experiments
@@
  • RQ1: How do the various synthesizers perform under the systematic assessment? What are the new findings?
@@
  • RQ3: How effective are the proposed metrics and evaluation framework?
Figure 2. Average ranking comparison for six HP synthesizers (outer means higher rank and better performance). Each vertex is the average performance rank of the method across 12 datasets, and each axis is the evaluation metric. “Fidelity($\mathcal{D}_{\text{train}}/\mathcal{D}_{\text{test}}$)” denotes the fidelity (i.e., Wasserstein distance) is evaluated on the training/test dataset. “MDS” is the proposed membership disclosure score (detailed in Section 4.2), and “MLA” is machine learning affinity (detailed in Section 5.2).
Figure 3. Visualization comparison of HP synthesizers on Bean dataset. Real data are in blue and synthetic data are in orange.
Figure 4. Average ranking comparison for four DP synthesizers. All methods offer provable privacy guarantees so we remove the privacy axis for comparison (the empirical privacy protections of these synthesizers are evaluated in Section 7.4).
Figure 5. Visualization comparison of DP synthesizers on Bean dataset. Real data are in blue and synthetic data are in orange.
7.1. Data Synthesis Models
We study a wide range of state-of-the-art data synthesizers, from statistical methods to deep generative models. We select them as they are either generally considered to perform best in practice (NIST, 2018, 2020), widely used (Xu et al., 2019; Papernot et al., 2018), or recently emerged (Kotelnikov et al., 2023; Borisov et al., 2023; Truda, 2023). These synthesizers can be categorized into two groups: heuristic private (HP) synthesizers and differentially private (DP) synthesizers.
@@ 7.1. Data Synthesis Models
  • CTGAN (Xu et al., 2019) is one of the most widely used HP synthesis algorithms. It utilizes generative adversarial networks to learn the tabular data distributions. Training techniques like conditional generation and Wasserstein loss (Gulrajani et al., 2017) are used.
  • TVAE (Xu et al., 2019) is the state-of-the-art variational autoencoder for tabular data synthesizer, which uses mode-specific normalization to tackle the non-Gaussian problems of continuous distributions.
  • TabDDPM (Kotelnikov et al., 2023) is the state-of-the-art diffusion model for data synthesis. It leverages the Gaussian diffusion process and the multinomial diffusion process to model continuous and discrete distributions respectively.
  • GReaT (Borisov et al., 2023) utilizes the large language model (LLM) for data synthesis. It converts records to textual representations for LLM and generates synthetic data with prompts.
@@
  • PGM (McKenna et al., 2021) is the state-of-the-art DP synthesizer, which uses probabilistic graphical models (McKenna et al., 2019) to learn the dependence of low-dimensional marginals for data synthesis. It won the NIST Differential Privacy Synthetic Data Challenge (NIST, 2018).
  • PrivSyn (Zhang et al., 2021) is a non-parametric DP synthesizer, which iteratively updates the synthetic dataset to make it match the target noise marginals. This method also shows strong performance in NIST competitions (NIST, 2018, 2020).
  • PATE-GAN (Jordon et al., 2018) shares a similar architecture with CTGAN, but leverages the Private Aggregation of Teacher Ensembles (PATE) mechanism (Papernot et al., 2018) to offer DP guarantees.
  • TableDiffusion (Truda, 2023) is a newly proposed diffusion model for tabular data synthesis, which uses Differentially Private Stochastic Gradient Descent (DP-SGD) to enforce privacy.
@@
Our goal is not to benchmark all possible synthesizers but to focus on the best-known and broad spectrum of state-of-the-art synthesizers. Hence, some synthesizers (Zhao et al., 2021; McKenna et al., 2022; Lee et al., 2023; Kim et al., 2022) are omitted as they are considered the variants of the algorithms we assess. Some approaches (e.g., TabDDPM and GReaT) are nearly simultaneously introduced, limiting opportunities for an extensive comparison. Our evaluation fills this gap by providing the first comprehensive evaluation of all state-of-the-art in a standardized setting.
Table 2. List of datasets used for evaluation.
@@ 7.2. Experimental Setups
Datasets. For systematic investigation of the performance of synthesis algorithms, we consider a diverse set of 12 real-world public datasets. These datasets have various sizes, nature, attributes, and distributions. We explicitly divide datasets into training and test with a ratio of 8:2, then split 20% of the training dataset as the validation set, which is used for model tuning. All datasets are publicly available and widely used, and the full list is detailed in Table 2.
@@
Evaluation Details. We use implementations from the original papers and make some small modifications to integrate them into our evaluation framework. For PGM and PrivSyn, we discretize numerical attributes into discrete ones by binning and setting the privacy budget to infinity when compared with HP synthesis algorithms. During the evaluation, we first tune the synthesizers with our proposed tuning objective using training and validation data and train them with optimal performance. Then, synthetic data are generated by trained generator $\mathsf{G}$ for evaluation, where we test 10 times and report the mean and standard deviation as the final score. The random seeds are fixed for each model to improve reproducibility. The complete hyperparameter search space of each synthesis algorithm is detailed in our technical report version.
@@
Table 3. Fidelity evaluation (i.e., Wasserstein distance, the lower score indicates the synthetic data is of higher quality) of synthesis algorithms on train data $\mathcal{D}_{\text{train}}$. The privacy budget $\epsilon$ of HP synthesizers is $\infty$ (the top part), and the budget for DP synthesizers is $1$ (the middle part). HALF and HISTOGRAM are the baselines that serve as the empirical upper bound and lower bound of the fidelity for HP synthesizers (detailed in Section 7.2). The best result is in bold.
@@
GReaT1 $0.050\scriptscriptstyle\pm\scriptstyle.002$ $0.049\scriptscriptstyle\pm\scriptstyle.003$ $0.076\scriptscriptstyle\pm\scriptstyle.002$ $0.037\scriptscriptstyle\pm\scriptstyle.003$ $0.050\scriptscriptstyle\pm\scriptstyle.006$ $0.020\scriptscriptstyle\pm\scriptstyle.001$ $0.055\scriptscriptstyle\pm\scriptstyle.005$ $0.055\scriptscriptstyle\pm\scriptstyle.003$ $0.022\scriptscriptstyle\pm\scriptstyle.005$ $0.094\scriptscriptstyle\pm\scriptstyle.004$ $0.019\scriptscriptstyle\pm\scriptstyle.001$
@@
HALF (upper bound) $020$002 $0.018\scriptscriptstyle\pm\scriptstyle.001$ $0.010\scriptscriptstyle\pm\scriptstyle.002$ $0.011\scriptscriptstyle\pm\scriptstyle.004$ $0.017\scriptscriptstyle\pm\scriptstyle.002$ $0.015\scriptscriptstyle\pm\scriptstyle.004$ $0.017\scriptscriptstyle\pm\scriptstyle.003$ $010$001 $0.012\scriptscriptstyle\pm\scriptstyle.004$ $0.009\scriptscriptstyle\pm\scriptstyle.001$ $0.026\scriptscriptstyle\pm\scriptstyle.004$ $0.006\scriptscriptstyle\pm\scriptstyle.000$
HISTOGRAM (lower bound) $0.213\scriptscriptstyle\pm\scriptstyle.013$ $0.1$003 $0.027\scriptscriptstyle\pm\scriptstyle.001$ $0.051\scriptscriptstyle\pm\scriptstyle.003$ $0.081\scriptscriptstyle\pm\scriptstyle.002$ $0.087\scriptscriptstyle\pm\scriptstyle.002$ $0.051\scriptscriptstyle\pm\scriptstyle.001$ $0.061\scriptscriptstyle\pm\scriptstyle.002$ $0.069\scriptscriptstyle\pm\scriptstyle.001$ $0.063\scriptscriptstyle\pm\scriptstyle.002$ $0.046\scriptscriptstyle\pm\scriptstyle.002$ $0.068\scriptscriptstyle\pm\scriptstyle.000$
@@
Table 4. Privacy evaluation (i.e., membership disclosure score, lower value means better empirical privacy protection) of HP synthesizers ($\epsilon=\infty$). SELF is the baseline that serves as the empirical lower bound of MDS (the upper bound of MDS is 0 by definition). The best result is in bold.